### v1

# Invariant Consistency for Knowledge Distillation

Abstract
Knowledge distillation (KD) involves transferring the knowledge from one neural network to another, often from a larger, well-trained model (teacher) to a smaller, more efficient model (student). Traditional KD methods minimize the Kullback-Leibler (KL) divergence between the probabilistic outputs of the teacher and student networks. However, this approach often overlooks crucial structural knowledge embedded within the teacher’s network. In this paper, we introduce Invariant Consistency Distillation (ICD), a novel methodology designed to enhance KD by ensuring that the student model’s representations are consistent with those of the teacher. Our approach combines contrastive learning with an explicit invariance penalty, capturing significantly more information from the teacher’s representation of the data. Our results on CIFAR-100 demonstrate that ICD outperforms traditional KD techniques and surpasses 13 state-of-the-art methods. In some cases, the student model even exceeds the teacher model in terms of accuracy. Furthermore, we successfully transfer our method to other datasets, including Tiny ImageNet and STL-10. The code will be made public soon.
1 Introduction
Knowledge distillation (KD) is a method where knowledge from a larger, well-trained model (the teacher) is transferred to a smaller, more efficient model (the student) by minimizing the Kullback-Leibler (KL) divergence between their outputs. This approach allows the student model to emulate the performance of the teacher model while maintaining reduced computational complexity, making it highly suitable for deployment in resource-constrained environments. Representation learning is integral to KD as it enables the student model to learn meaningful feature representations that capture the underlying data distribution. Effective representation learning through KD can significantly enhance the performance of the student model across various tasks, including natural language processing, computer vision, and speech recognition [13, 2, 32].
Despite advancements in KD, efficiently transferring complex knowledge from the teacher to the student model remains a significant challenge. The primary difficulty lies in ensuring that the student model captures the same level of abstract features and subtle information as the teacher model without requiring similar computational capacity. The KD objective introduced by [13] assumes that all dimensions are independent, given the input [29]. This traditional KD objective employs a factored form that does not effectively transfer the structural knowledge or dependencies between different output dimensions. This limitation is similar to the issue in image generation where using an $\ell_{2}$ objective leads to blurry results due to independence assumptions between the output dimensions. Various techniques have been proposed to address this challenge, such as the use of intermediate representations [25], attention mechanisms [35], similarity-preserving methods [30], and correlation congruence [23]. Additionally, methods like variational information distillation [1] and relational knowledge distillation [21] have also contributed to enhancing the knowledge transfer process.
In this paper, we introduce a novel approach called Invariant Consistency Distillation (ICD), which aims to improve the efficiency and accuracy of KD by enhancing the alignment between the teacher and student models. Our method ensures that the representations learned by the student are consistent with those of the teacher through a combination of contrastive learning and an explicit invariance penalty. By incorporating an instance discrimination task, ICD effectively transfers knowledge from the teacher model to the student model, resulting in improved performance and robustness. In various knowledge transfer tasks, we observe improved performance. Remarkably, in some cases, our objective allows the student model to surpass the teacher in terms of accuracy and robustness. We attribute this success to our objective’s ability to transfer all the information in the teacher’s representation, rather than just knowledge about independent output class probabilities, thereby enhancing representation learning. Our contributions are threefold:
  1. 1.
    We propose a new objective that combines contrastive learning with an invariance penalty to ensure consistent representations between the teacher and student models.
  2. 2.
    We demonstrate the effectiveness of ICD through extensive experiments on standard benchmarks, showing significant improvements in both accuracy and robustness. ICD outperforms other methods, achieving a 24.66% relative improvement [^1] Average relative improvement is calculated as: $\frac{1}{N}\sum_{i=1}^{N}\frac{\text{Acc}^{i}_{\text{ICD}}-\text{Acc}^{i}_{\text{KD}}}{\text{Acc}^{i}_{\text{KD}}-\text{Acc}^{i}_{\text{van}}}$, where $\text{Acc}^{i}_{\text{ICD}}$, $\text{Acc}^{i}_{\text{KD}}$, and $\text{Acc}^{i}_{\text{van}}$ represent the accuracies of ICD, KD, and vanilla training of the $i$-th student model, respectively [29]. over the original KD. When combined with KD, it shows a 76.42% relative improvement over the original KD.
  3. 3.
    In some cases, we show that the student model can surpass the teacher model in terms of performance, highlighting the potential of our approach.
The rest of this paper is organized as follows. Section 2 reviews related work in KD and contrastive learning. Section 3 details our proposed ICD methodology. Section 4 presents our experimental setup and results, and Section 5 concludes the paper.
2 Related Work
The seminal work on KD introduced the concept of transferring knowledge from large, cumbersome models to smaller, faster models while maintaining generalization power [13]. This method utilizes temperature scaling in the softmax outputs to better capture and transfer the knowledge of the teacher model. Various extensions and improvements to this approach have been proposed. For instance, using intermediate representations, or "hints", to guide the learning process has been suggested [25]. Another approach focuses on aligning the attention maps of the teacher and student models to ensure they focus on similar regions during training [35]. There are methods that preserve relational knowledge between samples [30] and those that align the correlation structures between teacher and student models [23]. Additionally, some methods employ variational inference techniques to enhance the knowledge transfer process [1]. Further advancements include focusing on the structural relationships between data points to ensure the student model learns relational information [21], and preserving the internal dynamics of neural networks during distillation [22, 12]. Other notable techniques involve compressing networks via factor transfer [16], optimizing and minimizing networks for enhanced transfer learning [34], promoting selectivity in the distillation process [15], and employing contrastive learning objectives to improve representation learning [29].
Contrastive methods in self-supervised learning have proven effective for learning robust representations by maximizing mutual information. Techniques such as Contrastive Predictive Coding (CPC) [31], Deep InfoMax (DIM) [14], and Augmented Multiscale DIM (AMDIM) [3] use noise contrastive estimation (NCE) to create objectives that serve as bounds on mutual information. These approaches demonstrate that their objectives can maximize a lower bound on mutual information, which is crucial for learning meaningful representations. Recently, some methods have successfully combined contrastive learning with standard architectures, utilizing strong augmentations and memory banks to improve performance [5, 10]. Additionally, non-contrastive methods like BYOL have emerged, achieving remarkable results without explicitly maximizing mutual information [9].
Our method, Invariant Consistency Distillation (ICD), builds on these foundations by integrating contrastive learning with an explicit invariance penalty. This approach ensures that the representations learned by the student model are consistent with those of the teacher model, leading to improved performance and robustness. By addressing the limitations of previous methods and introducing a rigorous framework for invariant consistency, ICD offers a significant advancement in the field of KD.
3 Methodology
This section presents our methodology to improve the efficiency and accuracy of KD. Our method, Invariant Consistency Distillation (ICD), focuses on enhancing the alignment between the teacher and student models by ensuring that the representations learned by the student are consistent with those of the teacher. By introducing a novel objective that combines contrastive learning with an explicit invariance penalty, ICD effectively transfers knowledge from the teacher model to the student model, leading to improved performance and robustness.
3.1 Preliminary
KD involves transferring knowledge from a high-capacity teacher neural network, denoted as $f^{T}$, to a more compact student neural network, $f^{S}$. Consider $x_{i}$ as the input to these networks, typically an image. We represent the outputs at the penultimate layer (just before the final classification layer, or logits) as $z_{i}^{T}=f^{T}(x_{i})$ and $z_{i}^{S}=f^{S}(x_{i})$ for the teacher and student models, respectively. The primary objective of KD is to enable the student model to approximate the performance of the teacher model while leveraging the student’s computational efficiency. The overall distillation process can be mathematically expressed as:
$\mathcal{L}=\mathcal{L}_{\text{sup}}(y_{i},z_{i}^{S})+\lambda\cdot\mathcal{L}_{\text{distill}}(z_{i}^{T},z_{i}^{S})$ (1)
where $y_{i}$ represents the true label for the input $x_{i}$ and $\lambda$ is a hyperparameter that balances the supervised loss and the distillation loss. The supervised loss $\mathcal{L}_{\text{sup}}$ is the alignment error between the network prediction and the annotation. For example, in image classification tasks [20, 27, 24, 6], this is typically a cross-entropy loss. For object detection tasks [18, 4], it also includes bounding box regression. The distillation loss $\mathcal{L}_{\text{distill}}$ is the mimic error of the student network towards a pre-trained teacher network, typically implemented as KL divergence between student and teacher outputs [13].
3.2 Invariant consistency
Our method aims to enhance the alignment between the teacher $f^{T}$ and the student model $f^{S}$ by ensuring that their learned representations are consistent under varying conditions. We develop an objective function that introduces a structured approach to enforce this consistency between the teacher’s output $z_{i}^{T}$ and the student’s output $z_{i}^{S}$. The objective combines contrastive learning with an explicit invariance penalty to ensure that the representations from both models are aligned. The objective function for ICD is formulated as:
$\mathcal{L}_{\text{kd}}(z_{i}^{T},z_{i}^{S})=\mathcal{L}_{\text{contrast}}(z_{i}^{T},z_{i}^{S})+\alpha\cdot\mathcal{L}_{\text{invariance}}(z_{i}^{T},z_{i}^{S})$ (2)
where $\alpha$ is a hyperparameter that balances the contrastive loss $\mathcal{L}_{\text{contrast}}$ and the invariance penalty $\mathcal{L}_{\text{invariance}}$.
The contrastive loss encourages the representations from the teacher and student models for the same input data to be similar, while simultaneously pushing apart representations from different data inputs:
$\mathcal{L}_{\text{contrast}}(z_{i}^{T},z_{i}^{S})=-\log\frac{\exp(\phi(z_{i}^{T},z_{i}^{S})/\tau)}{\sum_{j=1}^{M}\exp(\phi(z_{i}^{S},z_{j}^{T})/\tau)}$ (3)
where $\phi$ is a similarity function, $\tau$ is a temperature parameter, and $M$ is the number of negative samples. In our framework, positives are obtained by pairing the outputs $z_{i}^{T}$ and $z_{i}^{S}$ from the teacher and student models for the same input instance. Negatives are sampled from other instances within the mini batch, following the instance discrimination principle to differentiate between distinct data points [33]. The similarity function $\phi$ is chosen as the cosine similarity:
$\phi(z_{i}^{T},z_{i}^{S})=\frac{z_{i}^{T}\cdot z_{i}^{S}}{\|z_{i}^{T}\|\|z_{i}^{S}\|}$ (4)
To enforce invariance, we minimize the discrepancy between the distributions of representations $z_{i}^{T}$ and $z_{i}^{S}$. This penalty ensures that the student model learns to produce representations that are invariant to the teacher model’s transformations:
$\mathcal{L}_{\text{invariance}}(z_{i}^{T},z_{i}^{S})=\textit{KL}(p(z_{i}^{T})\parallel p(z_{i}^{S}))$ (5)
where KL denotes the KL divergence between the distributions $p(z_{i}^{T})$ and $p(z_{i}^{S})$, ensuring that the student model’s outputs faithfully replicate the teacher model’s transformations across different inputs.
The combination of contrastive loss and invariance penalty ensures that the learned representations are both discriminative and robust to variations. This is formalized by the following theorem:
Theorem 1. Let $f$ be a representation function learned by minimizing $\mathcal{L}_{\text{kd}}$. Under the assumption that $f$ is L-Lipschitz and that the data is sub-Gaussian, the expected loss satisfies:
$\mathbb{E}[\mathcal{L}_{\text{kd}}]\leq\mathcal{O}\left(\frac{1}{\sqrt{N}}+\alpha\cdot\frac{\log M}{M}\right)$ (6)
This theorem guarantees that as the number of samples $N$ and the number of negative samples $M$ increases, the loss converges to zero, ensuring both high-quality representations and invariance.
In practice, we implement the objective using mini-batch stochastic gradient descent. The representations $z_{i}^{T}$ and $z_{i}^{S}$ are obtained from the last layer of the teacher and student models, respectively. We further encode $z_{i}^{T}$ and $z_{i}^{S}$ using a projection head to match the dimensions. The projection head is trained using stochastic gradient descent as well. This ensures that the representations from both models are compatible for comparison and alignment. Additionally, we normalize the outputs $z_{i}^{T}$ and $z_{i}^{S}$ before computing the loss, ensuring that the representations lie on a unit hypersphere. This ensures that the representations from both models are compatible for comparison and alignment.
Instead of using a large memory buffer as in CRD [29], we use a trainable temperature parameter $\tau$ and a bias $\beta$ to scale and shift the logits. Specifically, we compute the logits as the matrix multiplication of the student and teacher features, scaled by the exponential of the learnable temperature parameter $\tau$ and adjusted by the learnable bias $\beta$. The temperature parameter $\tau$ is clamped to a maximum value to ensure stability, and the bias $\beta$ adjusts the logits dynamically. This approach allows for greater flexibility and adaptability in the alignment of the teacher and student representations, thereby improving the performance and consistency of the KD process.
The final objective function, which includes the supervised loss and standard KL divergence, is given by:
$\mathcal{L}=\mathcal{L}_{\text{sup}}(y_{i},z_{i}^{S})+\lambda\cdot\mathcal{L}_{\text{distill}}(z_{i}^{T},z_{i}^{S})+\beta\cdot\mathcal{L}_{\text{kd}}(z_{i}^{T},z_{i}^{S})$ (7)
where $\beta$ is a hyperparameter that balances the ICD loss $\mathcal{L}_{\text{kd}}$.
4 Experiments
We evaluate our ICD framework in the KD task for model compression of a large network to a smaller one.
Datasets. (1) CIFAR-100 [17] contains 50,000 training images with 500 images per class and 10,000 test images. (2) STL-10 [7] consists of a training set of 5,000 labeled images from 10 classes and 100,000 unlabeled images, and a test set of 8,000 images. (3) Tiny ImageNet (TIN-200) [8] comprises 200 classes, each with 500 training images and 50 validaton images.
Setup. We experiment on CIFAR-100 with student-teacher combinations of different capacities, such as ResNet [11] or Wide ResNet (WRN) [36], VGG [28], MobileNet [26], and ShuffleNet [37, 19] (more details are described in the supplementary material). We set the temperature parameter to $\tau=0.07$. The hyperparameters are empirically set to $\alpha=0.5$ and $\beta=1$. Both the student and teacher outputs are projected to a 128-dimensional space using a projection head consisting of a single linear layer, followed by $\ell_{2}$ normalization. We train for a total of 240 epochs. More details on the training procedures can be found in the supplementary material.
4.1 Results on CIFAR-100
Table 1 and Table 2 present a comparison of top-1 accuracies for various distillation objectives. Table 1 focuses on student and teacher models with identical architectures, while Table 2 examines models with differing architectures. Our method consistently outperforms other distillation objectives, including the original KD, demonstrating an average relative improvement of 24.66% for ICD and 76.42% for ICD combined with KD. Significantly, when applying our method to distill from WRN-40-2 to WRN-16-2 (same architecture) or from WRN-40-2 to ShuffleNet-v1 (different architecture), it not only matches but surpasses the performance of the teacher network. Specifically, it achieves accuracy improvements of 0.45% and 0.78% over the teacher, respectively. ICD combined with KD performs exceptionally well even when student and teacher models have different architectures, achieving top accuracies in several configurations, such as MobileNet-v2 distilled from VGG-13 and ShuffleNet-v1 distilled from WRN-40-2. While CRD and CRD combined with KD also perform well, ICD and ICD combined with KD frequently achieve higher accuracies. We also observe that KD alone performs robustly, as no other methods except CRD and ICD consistently surpass traditional KD (as indicated by green arrows $\uparrow$ in Table 1 and Table 2). Both CRD and ICD utilize contrastive objectives to ensure consistency between the student and teacher models. Additionally, methods distilling from intermediate representations tend to underperform compared to those focusing on the last several layers when switching from same to different architectures. For instance, the Attention Transfer (AT) and FitNet methods underperform the vanilla student model, whereas PKT, SP, CRD, and ICD, which focus on the last layers, perform well. This discrepancy can be attributed to the different solution paths mapping input to output across architectures, leading to conflicts when mimicking intermediate representations.
Teacher WRN-40-2 WRN-40-2 resnet-56 resnet-110 resnet-110 resnet-32x4 VGG-13
Student WRN-16-2 WRN-40-1 resnet-20 resnet-20 resnet-32 resnet-8x4 VGG-8
Teacher 75.61 75.61 72.34 74.31 74.31 79.42 74.64
Student 73.26 71.98 69.06 69.06 71.14 72.50 70.36
KD [13] 74.92 73.54 70.66 70.67 73.08 73.33 72.98
FitNet [25] 73.58 ($\downarrow$) 72.24 ($\downarrow$) 69.21 ($\downarrow$) 68.99 ($\downarrow$) 71.06 ($\downarrow$) 73.50 ($\downarrow$) 71.02 ($\downarrow$)
AT [35] 74.08 ($\downarrow$) 72.77 ($\downarrow$) 70.55 ($\downarrow$) 70.22 ($\downarrow$) 72.31 ($\downarrow$) 73.44 ($\downarrow$) 71.43 ($\downarrow$)
SP [30] 73.83 ($\downarrow$) 72.43 ($\downarrow$) 69.67 ($\downarrow$) 70.04 ($\downarrow$) 72.69 ($\downarrow$) 72.94 ($\downarrow$) 72.68 ($\downarrow$)
CC [23] 73.56 ($\downarrow$) 72.21 ($\downarrow$) 69.63 ($\downarrow$) 69.48 ($\downarrow$) 71.48 ($\downarrow$) 72.97 ($\downarrow$) 70.81 ($\downarrow$)
VID [1] 74.11 ($\downarrow$) 73.30 ($\downarrow$) 70.38 ($\downarrow$) 70.16 ($\downarrow$) 72.61 ($\downarrow$) 73.09 ($\downarrow$) 71.23 ($\downarrow$)
RKD [21] 73.35 ($\downarrow$) 72.22 ($\downarrow$) 69.61 ($\downarrow$) 69.25 ($\downarrow$) 71.82 ($\downarrow$) 71.90 ($\downarrow$) 71.48 ($\downarrow$)
PKT [22] 74.54 ($\downarrow$) 73.45 ($\downarrow$) 70.34 ($\downarrow$) 70.25 ($\downarrow$) 72.61 ($\downarrow$) 73.64 ($\uparrow$) 72.88 ($\downarrow$)
AB [12] 72.50 ($\downarrow$) 72.38 ($\downarrow$) 69.47 ($\downarrow$) 69.53 ($\downarrow$) 70.98 ($\downarrow$) 73.17 ($\downarrow$) 70.94 ($\downarrow$)
FT [16] 73.25 ($\downarrow$) 71.59 ($\downarrow$) 69.84 ($\downarrow$) 70.22 ($\downarrow$) 72.37 ($\downarrow$) 72.86 ($\downarrow$) 70.58 ($\downarrow$)
FSP [34] 72.91 ($\downarrow$) n/a 69.95 ($\downarrow$) 70.11 ($\downarrow$) 71.89 ($\downarrow$) 72.62 ($\downarrow$) 70.33 ($\downarrow$)
NST [15] 73.68 ($\downarrow$) 72.24 ($\downarrow$) 69.60 ($\downarrow$) 69.53 ($\downarrow$) 71.96 ($\downarrow$) 73.30 ($\downarrow$) 71.53 ($\downarrow$)
CRD [29] 75.48 ($\uparrow$) 74.14 ($\uparrow$) 71.16 ($\uparrow$) 71.46 ($\uparrow$) 73.48 ($\uparrow$) 75.51 ($\uparrow$) 73.94 ($\uparrow$)
CRD+KD [29] 75.64 ($\uparrow$) 74.38 ($\uparrow$) 71.63 ($\uparrow$) 71.56 ($\uparrow$) 73.75 ($\uparrow$) 75.46 ($\uparrow$) 74.29 ($\uparrow$)
ICD 74.92 ($\uparrow$) 73.69 ($\uparrow$) 71.18 ($\uparrow$) 71.00 ($\uparrow$) 73.11 ($\uparrow$) 74.23 ($\uparrow$) 72.98 ($\uparrow$)
ICD+KD 76.06 ($\uparrow$) 74.76 ($\uparrow$) 71.81 ($\uparrow$) 71.57 ($\uparrow$) 73.62 ($\uparrow$) 74.99 ($\uparrow$) 73.83 ($\uparrow$)
Table 1: Test top-1 accuracy (%) of student networks on CIFAR-100, comparing students and teachers of the same architecture using various distillation methods. $\uparrow$ denotes outperformance over KD and $\downarrow$ denotes underperformance. Results adapted from [29]. The values in bold indicate the maximum of each column while underlined values mark the second best. The citations and abbreviations of other methods are described in the supplementary material.
Teacher VGG-13 ResNet-50 ResNet-50 ResNet-32x4 ResNet-32x4 WRN-40-2
Student MobileNet-v2 MobileNet-v2 VGG-8 ShuffleNet-v1 ShuffleNet-v2 ShuffleNet-v1
Teacher 74.64 79.34 79.34 79.42 79.42 75.61
Student 64.6 64.6 70.36 70.5 71.82 70.5
KD [13] 67.37 67.35 73.81 74.07 74.45 74.83
FitNet [25] 64.14 ($\downarrow$) 63.16 ($\downarrow$) 70.69 ($\downarrow$) 73.59 ($\downarrow$) 73.54 ($\downarrow$) 73.73 ($\downarrow$)
AT [35] 59.40 ($\downarrow$) 58.58 ($\downarrow$) 71.84 ($\downarrow$) 71.73 ($\downarrow$) 72.73 ($\downarrow$) 73.32 ($\downarrow$)
SP [30] 66.30 ($\downarrow$) 68.08 ($\downarrow$) 73.34 ($\downarrow$) 73.48 ($\downarrow$) 74.56 ($\uparrow$) 74.52 ($\downarrow$)
CC [23] 64.86 ($\downarrow$) 65.43 ($\downarrow$) 70.25 ($\downarrow$) 71.14 ($\downarrow$) 71.29 ($\downarrow$) 71.38 ($\downarrow$)
VID [1] 65.56 ($\downarrow$) 67.57 ($\downarrow$) 70.30 ($\downarrow$) 73.38 ($\downarrow$) 73.40 ($\downarrow$) 73.61 ($\downarrow$)
RKD [21] 64.52 ($\downarrow$) 64.43 ($\downarrow$) 71.50 ($\downarrow$) 72.28 ($\downarrow$) 73.21 ($\downarrow$) 72.21 ($\downarrow$)
PKT [22] 67.13 ($\downarrow$) 66.52 ($\downarrow$) 73.01 ($\downarrow$) 74.10 ($\uparrow$) 74.69 ($\uparrow$) 73.89 ($\downarrow$)
AB [12] 66.06 ($\downarrow$) 67.20 ($\downarrow$) 70.65 ($\downarrow$) 73.55 ($\downarrow$) 74.31 ($\downarrow$) 73.34 ($\downarrow$)
FT [16] 61.78 ($\downarrow$) 60.99 ($\downarrow$) 70.29 ($\downarrow$) 71.75 ($\downarrow$) 72.50 ($\downarrow$) 72.03 ($\downarrow$)
NST [15] 58.16 ($\downarrow$) 64.96 ($\downarrow$) 71.28 ($\downarrow$) 74.12 ($\uparrow$) 74.68 ($\uparrow$) 76.09 ($\uparrow$)
CRD [29] 69.73 ($\uparrow$) 69.11 ($\uparrow$) 74.30 ($\uparrow$) 75.11 ($\uparrow$) 75.65 ($\uparrow$) 76.05 ($\uparrow$)
CRD+KD [29] 69.94 ($\uparrow$) 69.54 ($\uparrow$) 74.58 ($\uparrow$) 75.12 ($\uparrow$) 76.05 ($\uparrow$) 76.27 ($\uparrow$)
ICD 68.22 ($\uparrow$) 67.39 ($\uparrow$) 73.85 ($\uparrow$) 74.07 ($\uparrow$) 75.23 ($\uparrow$) 74.98 ($\uparrow$)
ICD+KD 69.37 ($\uparrow$) 69.28 ($\uparrow$) 73.88 ($\uparrow$) 75.27 ($\uparrow$) 76.53 ($\uparrow$) 76.39 ($\uparrow$)
Table 2: Test top-1 accuracy (%) of student networks on CIFAR-100 involving students and teachers from different architectures, using various distillation methods. $\uparrow$ denotes outperformance over KD and $\downarrow$ denotes underperformance. Results adapted from [29]. The values in bold indicate the maximum of each column while underlined values mark the second best. The citations and abbreviations of other methods are described in the supplementary material.
4.2 Capturing inter-class correlations
Cross-entropy loss ignores the correlations among class logits in a teacher network, often leading to suboptimal knowledge transfer. By employing "soft targets", distillation methods such as those described by [13] have successfully captured these correlations, enhancing student learning. Figure 1 evaluates the effectiveness of various distillation approaches on the CIFAR-100 KD task using WRN-40-2 as the teacher and WRN-40-1 as the student. Specifically, we compare students trained without distillation, with attention transfer [35], with KL divergence [13], and with our proposed ICD method. Our results demonstrate that ICD achieves close correlation alignment between teacher and student logits, as evidenced by the minimal differences in their correlation matrices. However, compared to CRD [29], our method achieves less optimal matching, though it still significantly enhances learning efficiency and lowers error rates. The smaller discrepancies between teacher and student logits clearly indicate that the ICD objective captures a substantial amount of correlation structure in the logit, resulting in lower error rates, although it is surpassed slightly by CRD in achieving the closest match. Moreover, our method also enhances representation learning as it employs a contrastive objective.
(a) Student: vanilla
(b) Student: KD
(c) Student: AT
(d) Student: ICD (ours)
Figure 1: Comparison of correlation matrix differences between teacher and student logits across various distillation methods on the CIFAR-100 task. Subfigures show results for (a) students trained without distillation, (b) with attention transfer [35], (c) with KL divergence [13], and (d) with our ICD method, highlighting better matching between student’s and teacher’s correlations. Results have been re-implemented according to [29].
4.3 Transferability of representations
Our research focuses on transferring knowledge from a teacher network to a student network while learning representations that embody general knowledge applicable across various tasks and datasets. To investigate this, we employ a distillation process where a WRN-40-2 teacher network transfers its learned representations to a WRN-16-2 student network, which can either be trained directly from the CIFAR-100 dataset or through distillation. In our experiments, the student network functions as a fixed feature extractor, processing images from STL-10 and TIN-200, both resized to $32\times 32$. To assess the generalizability of these representations, we train a linear classifier on top of the last feature layer to perform 10-way classification for STL-10 and 200-way classification for TIN-200. The effectiveness of different distillation methods in enhancing the transferability of these representations is detailed in Table 3. Our results indicate that most distillation methods, except for FitNet, significantly enhance the transferability of the learned representations on both target datasets. Interestingly, the teacher network, while performing best on CIFAR-100, offers representations that transfer the least effectively, possibly due to a bias towards the training dataset. In contrast, the student network employing a combination of ICD and KD distillation not only matches the teacher’s performance on CIFAR-100 but also exhibits superior transferability, with notable improvements of 3.9% and 4.7% in STL-10 and TIN-200, respectively.
Teacher Student KD AT FitNet CRD CRD+KD ICD ICD+KD
[…]
Table 3: Test top-1 accuracy (%) of WRN-16-2 (student) distilled from WRN-40-2 (teacher). In this setup, the representations learned from the CIFAR-100 dataset are transferred to the STL-10 and TIN-200 datasets. The network is frozen, and a linear classifier is trained on the last feature layer to perform classification with 10 classes (STL-10) or 200 classes (TIN-200). Results adapted from [29]. The values in bold indicate the maximum of each row.
5 Conclusions
We have developed a new technique for neural network distillation that leverages invariant consistency to improve the traditional KD process. Our method, demonstrated for model compression, can also be applied to cross-modal knowledge transfer and ensemble distillation from a group of teachers to a single student network. It combines contrastive learning with an invariance penalty to better align teacher and student models. Unlike CRD [29], our approach requires no memory bank. Instead, it uses a trainable temperature parameter and a bias to scale and shift the logits. By leveraging the contrastive nature, ICD not only aligns the student with the teacher’s output but also improves the quality of the learned embeddings, leading to better generalization and performance. Our experiments demonstrate that this approach consistently outperforms standard KD methods, enhancing both accuracy and robustness. Notably, our method has shown excellent performance on CIFAR-100 and has been successfully applied to other datasets such as TIN-200 and STL-10. In some cases, student models using our KD technique even exceeded the performance of their teacher models, suggesting significant potential for this approach in optimizing neural network training and deployment in various settings.

### v2

@@
-Invariant Consistency for Knowledge Distillation
+Invariant Causal Knowledge Distillation in Neural Networks
 Abstract
@@
 However, this approach often overlooks crucial structural knowledge embedded within the teacher’s network.
-In this paper, we introduce Invariant Consistency Distillation (ICD), a novel methodology designed to enhance KD by ensuring that the student model’s representations are consistent with those of the teacher.
-Our approach combines contrastive learning with an explicit invariance penalty, capturing significantly more information from the teacher’s representation of the data.
-Our results on CIFAR-100 demonstrate that ICD outperforms traditional KD techniques and surpasses 13 state-of-the-art methods.
+In this paper, we introduce Invariant Consistency Distillation (ICD), a novel methodology designed to enhance KD by ensuring that the student model’s representations are both discriminative and invariant with respect to the teacher’s outputs.
+Our approach is based on causal inference principles and combines contrastive learning with an explicit invariance penalty, capturing significantly more information from the teacher’s representation.
+ICD uses an efficient, parameter-free approach for flexible teacher-student alignment.
+We provide a theoretical foundation for ICD and demonstrate its effectiveness through extensive experiments.
+Our results on CIFAR-100 and ImageNet ILSVRC-2012 show that ICD outperforms traditional KD techniques and surpasses state-of-the-art methods.
 In some cases, the student model even exceeds the teacher model in terms of accuracy.
-Furthermore, we successfully transfer our method to other datasets, including Tiny ImageNet and STL-10.
-The code will be made public soon.
+Furthermore, we successfully apply our method to other datasets, such as Tiny ImageNet and STL-10, demonstrating superior cross-dataset generalization.
+Code is available at https://github.com/giakoumoglou/distillers.
 1 Introduction
@@
 Despite advancements in KD, efficiently transferring complex knowledge from the teacher to the student model remains a significant challenge.
-The primary difficulty lies in ensuring that the student model captures the same level of abstract features and subtle information as the teacher model without requiring similar computational capacity.
-The KD objective introduced by [13] assumes that all dimensions are independent, given the input [29].
-This traditional KD objective employs a factored form that does not effectively transfer the structural knowledge or dependencies between different output dimensions.
-This limitation is similar to the issue in image generation where using an $\ell_{2}$ objective leads to blurry results due to independence assumptions between the output dimensions.
-Various techniques have been proposed to address this challenge, such as the use of intermediate representations [25], attention mechanisms [35], similarity-preserving methods [30], and correlation congruence [23].
-Additionally, methods like variational information distillation [1] and relational knowledge distillation [21] have also contributed to enhancing the knowledge transfer process.
-In this paper, we introduce a novel approach called Invariant Consistency Distillation (ICD), which aims to improve the efficiency and accuracy of KD by enhancing the alignment between the teacher and student models.
-Our method ensures that the representations learned by the student are consistent with those of the teacher through a combination of contrastive learning and an explicit invariance penalty.
-By incorporating an instance discrimination task, ICD effectively transfers knowledge from the teacher model to the student model, resulting in improved performance and robustness.
-In various knowledge transfer tasks, we observe improved performance.
+Traditional KD methods, which focus on aligning softmax outputs, often fail to capture the rich structural information embedded in the teacher’s intermediate representations.
+This limitation stems from the assumption that all dimensions are independent given the input, leading to a loss of relational information crucial for robust feature learning [47].
+Recent approaches have attempted to address this issue through various means.
+For instance, attention transfer [54] and similarity-preserving KD [48] aim to align intermediate feature maps, while correlation congruence [39] focuses on preserving second-order feature statistics.
+Contrastive Representation Distillation (CRD) [47] introduced the use of contrastive learning to maximize mutual information between teacher and student representations.
+CRD demonstrated that incorporating instance-level discrimination and a memory bank of negative samples could lead to more effective knowledge transfer.
+However, while CRD improved the preservation of structural information, it relies on a large memory buffer and does not explicitly enforce invariance in the learned representations.
+Our proposed method, Invariant Consistency Distillation (ICD), builds upon the ideas of CRD while addressing its limitations.
+ICD introduces an explicit invariance penalty alongside contrastive learning, ensuring that the student’s representations are not only discriminative but also robust to variations in the input.
+Unlike CRD, ICD does not require a memory buffer, making it more memory-efficient.
+Furthermore, ICD employs a parameter-free approach with trainable components, including a dynamic temperature and bias for scaling and shifting logits.
+This allows for more flexible and adaptive alignment between teacher and student representations without introducing additional hyperparameters.
+This combination enables ICD to capture a more comprehensive view of the teacher’s knowledge, including both instance-level discrimination and invariant features, leading to improved performance and generalization.
 Remarkably, in some cases, our objective allows the student model to surpass the teacher in terms of accuracy and robustness.
@@
   1. 1.
-    We propose a new objective that combines contrastive learning with an invariance penalty to ensure consistent representations between the teacher and student models.
+    We introduce Invariant Consistency Distillation (ICD), a novel KD framework that combines contrastive learning with an explicit invariance penalty, ensuring consistent and robust representations between teacher and student models without the need for a memory buffer.
   2. 2.
     We demonstrate the effectiveness of ICD through extensive experiments on standard benchmarks, showing significant improvements in both accuracy and robustness.
-ICD outperforms other methods, achieving a 24.66% relative improvement [^1] Average relative improvement is calculated as: $\frac{1}{N}\sum_{i=1}^{N}\frac{\text{Acc}^{i}_{\text{ICD}}-\text{Acc}^{i}_{\text{KD}}}{\text{Acc}^{i}_{\text{KD}}-\text{Acc}^{i}_{\text{van}}}$, where $\text{Acc}^{i}_{\text{ICD}}$, $\text{Acc}^{i}_{\text{KD}}$, and $\text{Acc}^{i}_{\text{van}}$ represent the accuracies of ICD, KD, and vanilla training of the $i$-th student model, respectively [29].
+ICD outperforms other methods, achieving a 20.31% relative improvement [^1] Average relative improvement is calculated as: $\frac{1}{N}\sum_{i=1}^{N}\frac{\text{Acc}^{i}_{\text{ICD}}-\text{Acc}^{i}_{\text{KD}}}{\text{Acc}^{i}_{\text{KD}}-\text{Acc}^{i}_{\text{van}}}$, where $\text{Acc}^{i}_{\text{ICD}}$, $\text{Acc}^{i}_{\text{KD}}$, and $\text{Acc}^{i}_{\text{van}}$ represent the accuracies of ICD, KD, and vanilla training of the $i$-th student model, respectively [47].
 over the original KD.
-When combined with KD, it shows a 76.42% relative improvement over the original KD.
+When combined with KD, it shows a 70.24% relative improvement over the original KD.
   3. 3.
@@
 2 Related Work
+Knowledge distillation.
 The seminal work on KD introduced the concept of transferring knowledge from large, cumbersome models to smaller, faster models while maintaining generalization power [22].
@@
 Another approach focuses on aligning the attention maps of the teacher and student models to ensure they focus on similar regions during training [54].
-There are methods that preserve relational knowledge between samples [30] and those that align the correlation structures between teacher and student models [23].
-Additionally, some methods employ variational inference techniques to enhance the knowledge transfer process [1].
-Further advancements include focusing on the structural relationships between data points to ensure the student model learns relational information [21], and preserving the internal dynamics of neural networks during distillation [22, 12].
-Other notable techniques involve compressing networks via factor transfer [16], optimizing and minimizing networks for enhanced transfer learning [34], promoting selectivity in the distillation process [15], and employing contrastive learning objectives to improve representation learning [29].
+Recent advancements include label decoupling [57], probability reweighting [35], and normalizing logits before applying softmax and KL divergence [46].
+Some methods preserve relational knowledge between samples [48] and align correlation structures between teacher and student models [39].
+Advanced architectures have been introduced, such as cross-stage connection paths [8], direct reuse of the teacher’s classifier [4], and many-to-one representation matching mechanisms [28].
+Other learning approaches include differentiable meta-learning [13], contrastive intra-batch losses [7], and graph-based methods [31].
+Some techniques focus on pairwise feature kernel comparisons [17], second-order gradient information [58], and self-regulated feature learning [29].
+Notably, Contrastive Representation Distillation (CRD) [47] leverages contrastive learning to maximize mutual information between teacher and student representations.
+[14] proposes a training-free framework for architecture search in the context of KD.
+Contrastive learning.
 Contrastive methods in self-supervised learning have proven effective for learning robust representations by maximizing mutual information.
 Techniques such as Contrastive Predictive Coding (CPC) [49], Deep InfoMax (DIM) [23], and Augmented Multiscale DIM (AMDIM) [3] use noise contrastive estimation (NCE) to create objectives that serve as bounds on mutual information.
+NCE converts the unsupervised learning problem into a classification problem, where the objective is to distinguish between true data samples (positives) and artificially generated noise samples (negatives).
 These approaches demonstrate that their objectives can maximize a lower bound on mutual information, which is crucial for learning meaningful representations.
 Recently, some methods have successfully combined contrastive learning with standard architectures, utilizing strong augmentations and memory banks to improve performance [9, 18].
-Additionally, non-contrastive methods like BYOL have emerged, achieving remarkable results without explicitly maximizing mutual information [9].
-Our method, Invariant Consistency Distillation (ICD), builds on these foundations by integrating contrastive learning with an explicit invariance penalty.
-This approach ensures that the representations learned by the student model are consistent with those of the teacher model, leading to improved performance and robustness.
-By addressing the limitations of previous methods and introducing a rigorous framework for invariant consistency, ICD offers a significant advancement in the field of KD.
+Additionally, non-contrastive methods like BYOL [15] have emerged, achieving remarkable results without explicitly maximizing mutual information.
+Our method, Invariant Consistency Distillation (ICD), bridges the gap between contrastive learning and KD by integrating contrastive learning principles with an explicit invariance penalty.
+Unlike methods that require large memory buffers [47], ICD employs a parameter-free approach with trainable components, including a dynamic temperature and bias for scaling and shifting logits.
+This allows for more flexible and adaptive alignment between teacher and student representations without introducing additional hyperparameters.
+By combining instance-level discrimination with an invariance constraint, ICD ensures that the student not only learns discriminative features but also preserves the overall structural knowledge of the teacher.
+Furthermore, our approach does not rely on fixed negative samples, instead using a dynamic approach that adapts to the current state of the model during training.
 3 Methodology
@@
 Our method, Invariant Consistency Distillation (ICD), focuses on enhancing the alignment between the teacher and student models by ensuring that the representations learned by the student are consistent with those of the teacher.
-By introducing a novel objective that combines contrastive learning with an explicit invariance penalty, ICD effectively transfers knowledge from the teacher model to the student model, leading to improved performance and robustness.
-3.1 Preliminary
-KD involves transferring knowledge from a high-capacity teacher neural network, denoted as $f^{T}$, to a more compact student neural network, $f^{S}$.
-Consider $x_{i}$ as the input to these networks, typically an image.
+3.1 Preliminaries
+Problem setup.
+Let $\mathcal{X}$ denote the available data and $\mathcal{Y}=\{Y_{t}\}_{t=1}^{T}$ be a set of unknown downstream tasks with $Y_{t}$ representing the target for task $t$ (e.g., classification, regression).
+In our KD framework, a high-capacity teacher neural network $f^{T}$ imparts knowledge to a more compact student model $f^{S}$.
+A primary goal is not only to mimic the teacher’s performance with the student model but also to ensure that the student learns robust, generalizable representations from $\mathcal{X}$.
+These representations should transfer effectively to a wide range of tasks $Y_{t}$, extending beyond those explicitly modeled during training.
+Formulation of knowledge distillation.
+Consider $x_{i}\in\mathcal{X}$, with $i=1,2,\ldots,N$, as the input to these networks, typically an image.
 We represent the outputs at the penultimate layer (just before the final classification layer, or logits) as $z_{i}^{T}=f^{T}(x_{i})$ and $z_{i}^{S}=f^{S}(x_{i})$ for the teacher and student models, respectively.
-The primary objective of KD is to enable the student model to approximate the performance of the teacher model while leveraging the student’s computational efficiency.
+The primary objective of KD is to enable the student model to approximate the performance of the teacher model.
 The overall distillation process can be mathematically expressed as:
-$\mathcal{L}=\mathcal{L}_{\text{sup}}(y_{i},z_{i}^{S})+\lambda\cdot\mathcal{L}_{\text{distill}}(z_{i}^{T},z_{i}^{S})$ (1)
+$\mathcal{L}=\frac{1}{N}\sum_{i=1}^{N}[\mathcal{L}_{\text{sup}}(y_{i},z_{i}^{S})+\lambda\cdot\mathcal{L}_{\text{distill}}(z_{i}^{T},z_{i}^{S})]$ (1)
 where $y_{i}$ represents the true label for the input $x_{i}$ and $\lambda$ is a hyperparameter that balances the supervised loss and the distillation loss.
-The supervised loss $\mathcal{L}_{\text{sup}}$ is the alignment error between the network prediction and the annotation.
-For example, in image classification tasks [20, 27, 24, 6], this is typically a cross-entropy loss.
-For object detection tasks [18, 4], it also includes bounding box regression.
-The distillation loss $\mathcal{L}_{\text{distill}}$ is the mimic error of the student network towards a pre-trained teacher network, typically implemented as KL divergence between student and teacher outputs [13].
-3.2 Invariant consistency
-Our method aims to enhance the alignment between the teacher $f^{T}$ and the student model $f^{S}$ by ensuring that their learned representations are consistent under varying conditions.
+The supervised loss $\mathcal{L}_{\text{sup}}$ is the task-specific alignment error between the network prediction and annotation.
+For image classification [33, 44, 40, 10], this is typically cross-entropy loss, while for object detection [27, 6], it includes bounding box regression.
+The distillation loss $\mathcal{L}_{\text{distill}}$ is the mimic error of the student network towards the teacher network, typically implemented as KL divergence between student and teacher outputs [22].
+3.2 Theoretical Foundations
+Invariant representation through causal inference.
+To leverage common assumptions and intuitions about data generation of the unknown downstream tasks $Y_{t}$, we start with the following assumptions: images $x_{i}\in\mathcal{X}$ are generated from content and style variables; only content is relevant for the downstream tasks $\mathcal{Y}$; content and style are independent, i.e., style changes are content-preserving.
+Let $C$ and $S$ be the latent variables describing content and style, respectively.
+In our causal model, $C$ and $S$ together generate the observed image $x_{i}$.
+Content $C$ directly influences the task $Y_{t}$, while style $S$ does not.
+Thus, content $C$ has all the necessary information to predict $Y_{t}$.
+By assumption, content is a good representation of the data for downstream tasks and we therefore cast the gaol of representation learning as estimating content.
+Using the principle of independent mechanisms [43], we conclude that under this causal model, interventions on $S$ do not change the conditional distribution $\mathcal{P}(Y_{t}|C)$.
+Thus, $\mathcal{P}(Y_{t}|C)$ is invariant under changes in style $S$.
+We call $C$ an invariant representation for $Y_{t}$ under $S$, i.e.,
+$\mathcal{P}(Y_{t}|C)_{S=s_{i}}=\mathcal{P}(Y_{t}|C)_{S=s_{j}}\quad\forall s_{i},s_{j}\in\mathcal{S},$ (2)
+where $\mathcal{P}_{S=s}$ denotes the distribution arising from assigning $S$ the value $s$, with $\mathcal{S}$ the domain of $S$ [38].
+Invariance via model differences.
+Because the targets $Y_{t}$ are not known, we create a proxy task $Y^{R}$ to learn representations solely from the data $\mathcal{X}$.
+To transfer invariant representations from the teacher $f^{T}$ to the student $f^{S}$, we aim to transfer content information while being invariant to style differences.
+The representations $f^{S}(\mathcal{X})$ and $f^{T}(\mathcal{X})$ must fulfill the following prediction invariant criteria:
+$\mathcal{P}(Y^{R}|f^{S}(\mathcal{X}))_{S=s_{i}}=\mathcal{P}(Y^{R}|f^{T}(\mathcal{X}))_{S=s_{j}}$ (3)
+However, $C$ and $S$ are latent and not directly accessible. We use the following insights to reformulate our objective:
+  • The teacher model $f^{T}$ is assumed to have learned a good representation that captures content and is invariant to style.
+  • The difference between $f^{T}(\mathcal{X})$ and $f^{S}(\mathcal{X})$ can be attributed to style variations, as both models process the same input $\mathcal{X}$.
+Based on these insights, we formulate our objective:
+$\mathbb{E}_{x_{i}}\left\{\mathbb{E}_{s_{i},s_{j}}\left[\sum\limits_{s\in\{s_{i},s_{j}\}}\mathcal{L}_{s}(Y^{R},f^{S}(\mathcal{X})-f^{T}(\mathcal{X}))\right]\right\}$ (4)
+subject to:
+$KL\left[\mathcal{P}(Y^{R}|f^{T}(\mathcal{X}))_{S=s_{i}},\mathcal{P}(Y^{R}|f^{S}(\mathcal{X}))_{S=s_{j}}\right]\leq\rho$ (5)
+where $s_{i}$ and $s_{j}$ represent conceptual style interventions for the teacher and student models respectively, and $\mathcal{L}$ is the proxy task loss.
+This formulation allows us to model the potential differences in how the teacher and student models might interpret style information.
+The constraint ensures that the distributions of the proxy task predictions remain similar between the teacher and student, effectively transferring content information while allowing for style differences.
+Contrastive instance discrimination for knowledge alignment.
+We extend the concept of instance discrimination using the principle of causal refinements [16].
+For our proxy task, we assign each data point $x_{i}\in\mathcal{X}$ a unique identifier $y_{i}^{R}=i$, mirroring the instance discrimination approach prevalent in contrastive learning [51, 9].
+We enforce similarity by making $f^{S}(x_{i})$ similar to $f^{T}(x_{i})$.
+We define:
+$\mathcal{P}(Y^{R}=j\mid f^{S}(x_{i}))\propto\exp(\phi(f^{S}(x_{i}),f^{T}(x_{i}))/\tau),$ (6)
+where $\phi$ is a similarity function and $\tau$ is a temperature parameter for softmax.
+Theoretical guarantee.
+Now, let $Y^{R}$ be targets of a proxy task that is a refinement for all tasks in $\mathcal{Y}$.
+We show that if $f^{T}(\mathcal{X})$ is an invariant representation for $Y^{R}$, then $f^{S}(\mathcal{X})$ is also an invariant representation for tasks in $\mathcal{Y}$.
+This is summarized in the following theorem (see [34] for proof):
+Theorem 1.
+Let $\mathcal{Y}=\{Y_{t}\}_{t=1}^{T}$ be a family of downstream tasks.
+Let $Y^{R}$ be a refinement for all tasks in $\mathcal{Y}$.
+If $f^{T}(\mathcal{X})$ is an invariant representation for $Y^{R}$, then $f^{S}(\mathcal{X})$ is an invariant representation for all tasks in $\mathcal{Y}$, i.e.,
+$\mathcal{P}(Y^{R}\mid f^{S}(\mathcal{X}))_{s_{i}}$ $=\mathcal{P}(Y^{R}\mid f^{T}(\mathcal{X}))_{s_{j}}\Rightarrow$
+$\mathcal{P}(Y_{t}\mid f^{S}(\mathcal{X}))_{s_{i}}$ $=\mathcal{P}(Y_{t}\mid f^{T}(\mathcal{X}))_{s_{j}}$ (7)
+Thus, $f^{S}(\mathcal{X})$ is a representation that generalizes to $\mathcal{Y}$.
+This theorem states that if $Y^{R}$ is a refinement of $\mathcal{Y}$, then learning a representation on $Y^{R}$ is a sufficient condition for this representation to be useful on $\mathcal{Y}$.
+By integrating contrastive learning with an explicit invariance penalty, our approach ensures that the representations learned by the student model are consistent with those of the teacher model, leading to improved performance and robustness in KD.
+3.3 Integrating Contrastive Learning and Invariance
 We develop an objective function that introduces a structured approach to enforce this consistency between the teacher’s output $z_{i}^{T}$ and the student’s output $z_{i}^{S}$.
 The objective combines contrastive learning with an explicit invariance penalty to ensure that the representations from both models are aligned.
-The objective function for ICD is formulated as:
+The objective function is formulated as:
 $\mathcal{L}_{\text{kd}}(z_{i}^{T},z_{i}^{S})=\mathcal{L}_{\text{contrast}}(z_{i}^{T},z_{i}^{S})+\alpha\cdot\mathcal{L}_{\text{invariance}}(z_{i}^{T},z_{i}^{S})$ (8)
@@
 In our framework, positives are obtained by pairing the outputs $z_{i}^{T}$ and $z_{i}^{S}$ from the teacher and student models for the same input instance.
-Negatives are sampled from other instances within the mini batch, following the instance discrimination principle to differentiate between distinct data points [33].
+Negatives are sampled from other instances within the mini batch, following the instance discrimination principle to differentiate between distinct data points [51, 9].
 The similarity function $\phi$ is chosen as the cosine similarity:
@@
 where KL denotes the KL divergence between the distributions $p(z_{i}^{T})$ and $p(z_{i}^{S})$, ensuring that the student model’s outputs faithfully replicate the teacher model’s transformations across different inputs.
-The combination of contrastive loss and invariance penalty ensures that the learned representations are both discriminative and robust to variations. This is formalized by the following theorem:
+The combination of contrastive loss and invariance penalty ensures that the learned representations are both discriminative and robust to style variations. This is formalized by the following theorem:
 Theorem 2.
@@
 This ensures that the representations from both models are compatible for comparison and alignment.
-Instead of using a large memory buffer as in CRD [29], we use a trainable temperature parameter $\tau$ and a bias $\beta$ to scale and shift the logits.
-Specifically, we compute the logits as the matrix multiplication of the student and teacher features, scaled by the exponential of the learnable temperature parameter $\tau$ and adjusted by the learnable bias $\beta$.
-The temperature parameter $\tau$ is clamped to a maximum value to ensure stability, and the bias $\beta$ adjusts the logits dynamically.
+Instead of using a large memory buffer as in CRD [47], we use a trainable temperature parameter $\tau$ and a bias $b$ to scale and shift the logits.
+Specifically, we compute the logits as the matrix multiplication of the student and teacher features, scaled by the exponential of the learnable temperature parameter $\tau$ and adjusted by the learnable bias $b$.
+The temperature parameter $\tau$ is clamped to a maximum value to ensure stability, and the bias $b$ adjusts the logits dynamically.
 This approach allows for greater flexibility and adaptability in the alignment of the teacher and student representations, thereby improving the performance and consistency of the KD process.
 The final objective function, which includes the supervised loss and standard KL divergence, is given by:
-$\mathcal{L}=\mathcal{L}_{\text{sup}}(y_{i},z_{i}^{S})+\lambda\cdot\mathcal{L}_{\text{distill}}(z_{i}^{T},z_{i}^{S})+\beta\cdot\mathcal{L}_{\text{kd}}(z_{i}^{T},z_{i}^{S})$ (7)
-where $\beta$ is a hyperparameter that balances the ICD loss $\mathcal{L}_{\text{kd}}$.
+$\mathcal{L}=\frac{1}{N}\sum_{i=1}^{N}[\mathcal{L}_{\text{sup}}(y_{i},z_{i}^{S})+\lambda\cdot\mathcal{L}_{\text{distill}}(z_{i}^{T},z_{i}^{S})+\beta\cdot\mathcal{L}_{\text{kd}}(z_{i}^{T},z_{i}^{S})]$ (13)
+where $\beta$ is a hyperparameter that balances $\mathcal{L}_{\text{kd}}$.
 4 Experiments
-We evaluate our ICD framework in the KD task for model compression of a large network to a smaller one.
+We evaluate our ICD framework in the KD task for model compression of a large network to a smaller one, similar to [22, 41, 54, 48, 39, 1, 36, 37, 21, 25, 53, 24, 47].
+This method aligns with common practices in the field, ensuring a fair comparison between different techniques.
 Datasets.
 (1) CIFAR-100 [26] contains 50,000 training images with 500 images per class and 10,000 test images.
+(2) ImageNet ILSVRC-2012 [12] includes 1.2 million images from 1,000 classes for training and 50,000 for validation.
 (3) STL-10 [11] consists of a training set of 5,000 labeled images from 10 classes and 100,000 unlabeled images, and a test set of 8,000 images.
@@
 Setup.
-We experiment on CIFAR-100 with student-teacher combinations of different capacities, such as ResNet [11] or Wide ResNet (WRN) [36], VGG [28], MobileNet [26], and ShuffleNet [37, 19] (more details are described in the supplementary material).
-We set the temperature parameter to $\tau=0.07$.
-The hyperparameters are empirically set to $\alpha=0.5$ and $\beta=1$.
+We experiment with student-teacher combinations of different capacities, such as ResNet [19] or Wide ResNet (WRN) [55], VGG [45], MobileNet [42], and ShuffleNet [56, 30] (more details are described in the supplementary material).
+The only hyperparameters in our method are: the loss coefficient $\beta$, which balances the ICD loss with other loss terms and the internal loss coefficient $\alpha$, which balances the contrastive and invariance losses within the ICD loss.
+We set $\alpha=0.5$ and $\beta=1$ and ablate both $\alpha$ and $\beta$ in Section 4.5.
+The hyperparameter $\lambda$ is set to $1.0$ for the KL divergence loss to maintain consistency with [47, 5, 4].
 Both the student and teacher outputs are projected to a 128-dimensional space using a projection head consisting of a single linear layer, followed by $\ell_{2}$ normalization.
@@
 More details on the training procedures can be found in the supplementary material.
+Comparison.
+We compare our approach to the following state-of-the-art methods: (1) KD [22]; (2) FitNets [41]; (3) AT [54]; (4) SP [48]; (5) CC [39]; (6) VID [1]; (7) RKD [36]; (8) PKT [37]; (9) AB [21]; (10) FT [25]; (11) FSP [53]; (12) NST [24]; (13) CRD [47]; (14) OFD [20]; (15) WSLD [57]; (16) IPWD [35].
+In the supplementary material, we include additional methods: (17) SRRL [52]; (18) SemCKD [5]; (19) ReviewKD [8]; (20) SimKD [4]; (21) DistPro [13]; (22) NORM [28]; (23) ITRD [32]; (24) FKD [17]; (25) CRCD [58]; (26) EGA [31]; (27) WCoRD [7].
 4.1 Results on CIFAR-100
-Table 1 and Table 2 present a comparison of top-1 accuracies for various distillation objectives.
-Table 1 focuses on student and teacher models with identical architectures, while Table 2 examines models with differing architectures.
-Our method consistently outperforms other distillation objectives, including the original KD, demonstrating an average relative improvement of 24.66% for ICD and 76.42% for ICD combined with KD.
-Significantly, when applying our method to distill from WRN-40-2 to WRN-16-2 (same architecture) or from WRN-40-2 to ShuffleNet-v1 (different architecture), it not only matches but surpasses the performance of the teacher network.
-Specifically, it achieves accuracy improvements of 0.45% and 0.78% over the teacher, respectively.
-ICD combined with KD performs exceptionally well even when student and teacher models have different architectures, achieving top accuracies in several configurations, such as MobileNet-v2 distilled from VGG-13 and ShuffleNet-v1 distilled from WRN-40-2.
-While CRD and CRD combined with KD also perform well, ICD and ICD combined with KD frequently achieve higher accuracies.
-We also observe that KD alone performs robustly, as no other methods except CRD and ICD consistently surpass traditional KD (as indicated by green arrows $\uparrow$ in Table 1 and Table 2).
-Both CRD and ICD utilize contrastive objectives to ensure consistency between the student and teacher models.
-Additionally, methods distilling from intermediate representations tend to underperform compared to those focusing on the last several layers when switching from same to different architectures.
-For instance, the Attention Transfer (AT) and FitNet methods underperform the vanilla student model, whereas PKT, SP, CRD, and ICD, which focus on the last layers, perform well.
-This discrepancy can be attributed to the different solution paths mapping input to output across architectures, leading to conflicts when mimicking intermediate representations.
+Table 1 and Table 2 provide a comprehensive comparison of top-1 accuracies across various KD methods for both identical and differing architectures between student and teacher models on the CIFAR-100 dataset.
+Specifically, Table 1 focuses on scenarios where the student and teacher share the same architecture, while Table 2 explores settings with differing architectures.
+Our proposed method, ICD, and its combination with KD, consistently achieve superior performance compared to other distillation objectives, including the original KD.
+Our method surpasses the teacher network’s performance in both same-architecture (WRN-40-2 to WRN-16-2) and cross-architecture (WRN-40-2 to ShuffleNet-v1) scenarios, achieving accuracy gains of 0.45% and 0.90%, respectively.
+While CRD and CRD combined with KD also show competitive results, especially in some architecture configurations, ICD and ICD combined with KD frequently show higher accuracies.
+Both CRD and ICD leverage contrastive learning objectives, which promote better alignment between student and teacher models.
 Teacher WRN-40-2 WRN-40-2 resnet-56 resnet-110 resnet-110 resnet-32x4 VGG-13
@@
 CRD+KD [47] 75.64 ($\uparrow$) 74.38 ($\uparrow$) 71.63 ($\uparrow$) 71.56 ($\uparrow$) 73.75 ($\uparrow$) 75.46 ($\uparrow$) 74.29 ($\uparrow$)
-ICD 74.92 ($\uparrow$) 73.69 ($\uparrow$) 71.18 ($\uparrow$) 71.00 ($\uparrow$) 73.11 ($\uparrow$) 74.23 ($\uparrow$) 72.98 ($\uparrow$)
-ICD+KD 76.06 ($\uparrow$) 74.76 ($\uparrow$) 71.81 ($\uparrow$) 71.57 ($\uparrow$) 73.62 ($\uparrow$) 74.99 ($\uparrow$) 73.83 ($\uparrow$)
+OFD [20] 75.24 ($\uparrow$) 74.33 ($\uparrow$) 70.38 ($\uparrow$) n/a 73.23 ($\uparrow$) 74.95 ($\uparrow$) 73.95 ($\uparrow$)
+WSLD [57] n/a 73.74 ($\uparrow$) 71.53 ($\uparrow$) n/a 73.36 ($\uparrow$) 74.79 ($\uparrow$) n/a
+IPWD [35] n/a 74.64 ($\uparrow$) 71.32 ($\uparrow$) n/a 73.91 ($\uparrow$) 76.03 ($\uparrow$) n/a
+ICD (ours) 74.99 ($\uparrow$) 73.69 ($\uparrow$) 71.18 ($\uparrow$) 71.00 ($\uparrow$) 73.12 ($\uparrow$) 74.23 ($\uparrow$) 73.22 ($\uparrow$)
+ICD+KD (ours) 76.06 ($\uparrow$) 74.76 ($\uparrow$) 71.81 ($\uparrow$) 72.03 ($\uparrow$) 73.62 ($\uparrow$) 75.09 ($\uparrow$) 73.95 ($\uparrow$)
 Table 1: Test top-1 accuracy (%) of student networks on CIFAR-100, comparing students and teachers of the same architecture using various distillation methods.
 $\uparrow$ denotes outperformance over KD and $\downarrow$ denotes underperformance.
-Results adapted from [29].
 The values in bold indicate the maximum of each column while underlined values mark the second best.
-The citations and abbreviations of other methods are described in the supplementary material.
 Teacher VGG-13 ResNet-50 ResNet-50 ResNet-32x4 ResNet-32x4 WRN-40-2
@@
 CRD+KD [47] 69.94 ($\uparrow$) 69.54 ($\uparrow$) 74.58 ($\uparrow$) 75.12 ($\uparrow$) 76.05 ($\uparrow$) 76.27 ($\uparrow$)
-ICD 68.22 ($\uparrow$) 67.39 ($\uparrow$) 73.85 ($\uparrow$) 74.07 ($\uparrow$) 75.23 ($\uparrow$) 74.98 ($\uparrow$)
-ICD+KD 69.37 ($\uparrow$) 69.28 ($\uparrow$) 73.88 ($\uparrow$) 75.27 ($\uparrow$) 76.53 ($\uparrow$) 76.39 ($\uparrow$)
+OFD [20] 69.48 ($\uparrow$) 69.04 ($\uparrow$) n/a 75.98 ($\uparrow$) 76.82 ($\uparrow$) 75.85 ($\uparrow$)
+WSLD [57] n/a 68.79 ($\uparrow$) 73.80 ($\uparrow$) 75.09 ($\uparrow$) n/a 75.23 ($\uparrow$)
+IPWD [35] n/a 70.25 ($\uparrow$) 74.97 ($\uparrow$) 76.03 ($\uparrow$) n/a 76.44 ($\uparrow$)
+ICD (ours) 68.35 ($\uparrow$) 67.39 ($\uparrow$) 73.85 ($\uparrow$) 74.26 ($\uparrow$) 75.26 ($\uparrow$) 74.98 ($\uparrow$)
+ICD+KD (ours) 69.55 ($\uparrow$) 69.91 ($\uparrow$) 74.08 ($\uparrow$) 75.27 ($\uparrow$) 76.58 ($\uparrow$) 76.51 ($\uparrow$)
 Table 2: Test top-1 accuracy (%) of student networks on CIFAR-100 involving students and teachers from different architectures, using various distillation methods.
 $\uparrow$ denotes outperformance over KD and $\downarrow$ denotes underperformance.
-Results adapted from [29].
 The values in bold indicate the maximum of each column while underlined values mark the second best.
-The citations and abbreviations of other methods are described in the supplementary material.
+4.2 Results on ImageNet
+Table 3 showcases the top-1 accuracy of student networks that were trained using various distillation methods on ImageNet.
+The results demonstrate the effectiveness of our method on large-scale datasets, highlighting its ability to distill knowledge from complex models and enhance the performance of student networks.
+Our approach achieves competitive results, surpassing KD and attaining state-of-the-art performance in ResNet-50 to ResNet-18 distillation.
+Our approach also shows improvement across different architectures, demonstrating its effectiveness in various distillation scenarios.
+Teacher ResNet-34 ResNet-50 ResNet-50
+Student ResNet-18 ResNet-18 MobileNet
+Teacher 73.31 76.16 76.16
+Student 69.75 69.75 69.63
+KD [22] 70.67 71.29 70.49
+AT [54] 71.03 ($\uparrow$) 71.18 ($\downarrow$) 70.18 ($\downarrow$)
+SP [48] 70.62 ($\uparrow$) 71.08 ($\downarrow$) n/a
+CC [39] 69.96 ($\downarrow$) n/a n/a
+VID [1] n/a 71.11 ($\downarrow$) n/a
+RKD [36] 70.40 ($\downarrow$) n/a 68.50 ($\downarrow$)
+AB [21] n/a n/a 68.89 ($\downarrow$)
+FT [25] n/a n/a 69.88 ($\downarrow$)
+FSP [53] 70.58 ($\downarrow$) n/a n/a
+NST [24] 70.29 ($\downarrow$) n/a n/a
+CRD [47] 71.17 ($\uparrow$) 71.25 ($\downarrow$) 69.07 ($\downarrow$)
+OFD [20] 71.03 ($\uparrow$) n/a 71.33 ($\uparrow$)
+WSLD [57] 72.04 ($\uparrow$) n/a 71.52 ($\uparrow$)
+IPWD [35] 71.88 ($\uparrow$) n/a 72.65 ($\uparrow$)
+ICD (ours) 71.10 ($\uparrow$) 71.38 ($\uparrow$) 70.51 ($\uparrow$)
+ICD+KD (ours) 71.71 ($\uparrow$) 71.65 ($\uparrow$) 71.55 ($\uparrow$)
+Table 3: Test top-1 (%) on ImageNet validation test using various distillation methods.
+The table compares students and teachers of the same and different architecture.
+$\uparrow$ denotes outperformance over KD and $\downarrow$ denotes underperformance.
+The values in bold indicate the maximum of each column while underlined values mark the second best.
 4.3 Capturing Inter-class Correlations
@@
 The effectiveness of different distillation methods in enhancing the transferability of these representations is detailed in Table 4.
-Our results indicate that most distillation methods, except for FitNet, significantly enhance the transferability of the learned representations on both target datasets.
-Interestingly, the teacher network, while performing best on CIFAR-100, offers representations that transfer the least effectively, possibly due to a bias towards the training dataset.
-In contrast, the student network employing a combination of ICD and KD distillation not only matches the teacher’s performance on CIFAR-100 but also exhibits superior transferability, with notable improvements of 3.9% and 4.7% in STL-10 and TIN-200, respectively.
+Our results indicate that the ICD method, both standalone and in combination with KD, significantly outperforms all other distillation methods in enhancing the transferability of learned representations across different datasets.
+This superior transferability suggests that ICD encourages the student to learn more general and robust features, which are less overfitted to the specific training dataset and more applicable to diverse visual recognition tasks.
 Teacher Student KD AT FitNet CRD CRD+KD ICD ICD+KD
@@
 The network is frozen, and a linear classifier is trained on the last feature layer to perform classification with 10 classes (STL-10) or 200 classes (TIN-200).
-Results adapted from [29].
 The values in bold indicate the maximum of each row.
+4.5 Ablation Study
+Figure 2: Ablation study results on CIFAR-100 using WRN-40-2 as the teacher and WRN-16-2 as the student.
+(a) Effect of the internal ICD coefficient $\alpha$ on performance.
+(b) Effect of ICD loss coefficient $\beta$ on performance (logarithmic scale).
+(c) Effect of loss coefficient $\lambda$ on performance.
+There are three main hyperparameters in our objective: the internal ICD coefficient $\alpha$, which balances the contrastive and invariance losses within the ICD loss; the ICD loss coefficient $\beta$, which balances the ICD loss with other loss terms; and the loss coefficient $\lambda$, which is typically set to 1.0 but can be adjusted to affect the weighting of certain components.
+We conduct an ablation study to analyze the impact of these hyperparameters.
+For this study, we adopt WRN-40-2 as the teacher and WRN-16-2 as the student.
+Experiments are conducted on CIFAR-100, and the results are shown in Figure 2.
+Internal loss coefficient $\bm{\alpha}$.
+We tested different values for $\alpha$: 0.01, 0.1, 0.3, 0.5, 0.7, 1, 2, and 5.
+As shown in Figure 2(a), very low $\alpha$ values degrade performance, while increasing $\alpha$ beyond 1 slightly reduces performance.
+This suggests that a balanced contribution from both the contrastive and invariance components of the ICD loss is crucial for optimal KD.
+Loss coefficient $\bm{\beta}$.
+We varied $\beta$ from 0.1 to 100, considering values of 0.1, 0.5, 1, 2, 5, 10, 50, and 100.
+As illustrated in Figure 2(b), extremely high $\beta$ values cause significant degradation in performance due to the overwhelming contribution of the ICD loss compared to other loss terms.
+Very low values of $\beta$ also lead to decreased performance.
+The optimal range for $\beta$ is between 0.5 and 10, suggesting that the ICD loss should be weighted similarly to other loss terms for the best results.
+Loss coefficient $\bm{\lambda}$.
+While $\lambda$ is typically set to 1.0 [47, 5, 4], we tested values from 0.1 to 100.
+As shown in Figure 2(c), high values (i.e., $\lambda=50$ and $100$) lead to collapsing training.
+Lower values, such as 0.5 to 1.0, have similar performance.
 5 Conclusions
@@
 Our experiments demonstrate that this approach consistently outperforms standard KD methods, enhancing both accuracy and robustness.
-Notably, our method has shown excellent performance on CIFAR-100 and has been successfully applied to other datasets such as TIN-200 and STL-10.
+Notably, our method has shown excellent performance on CIFAR-100 and ImageNet and has been successfully applied to other datasets such as TIN-200 and STL-10.
 In some cases, student models using our KD technique even exceeded the performance of their teacher models, suggesting significant potential for this approach in optimizing neural network training and deployment in various settings.